OpenStudy (anonymous):

please help Show all work in dividing 4 t squared minus 1 all over 15 t plus 10 divided by 4 t squared minus 4 t plus 1 all over 3 t squared plus 5 t plus 2

OpenStudy (anonymous):

anyone?

OpenStudy (rajee_sam):

\[\frac{ \frac{ (4t ^{2} -1) }{ (15t + 10) } }{\frac{ (4t ^{2} -4t + 1) }{ (3t ^{2} + 5t +2) } }\] Is this your problem??

OpenStudy (rajee_sam):

@jannaruss26

OpenStudy (rajee_sam):

\[\frac{ 4t ^{2} - 1 }{ 15t + 10 } \times \frac{ 3t ^{2} + 5t + 2 }{4t ^{2}-4t + 1 }\] Now we have factor all the numerators and denominators \[\frac{ (2t +1 ) (2t -1) }{ 5 (3t + 2) } \times \frac{ (3t + 2) (t + 1) }{(2t -1)(2t - 1) }\]

OpenStudy (rajee_sam):

Now If you simplify the expression it will be \[\frac{ (2t + 1) ( t + 1) }{ 5 (2t - 1) }\]

OpenStudy (anonymous):

thank you!!!